OpenStudy (anonymous):

Really need help on this assignment thats due tonight at midnight eastern time.. Will give medals to whoever helps please

OpenStudy (anonymous):

Throw it at me

OpenStudy (anonymous):

Happy Birthday! Or maybe Happy Un–Birthday? Either way, here is a fun birthday trick. Take the month you were born in. Multiply it by 4, add 13, multiply by 25, subtract 200, add the day you were born, multiply by 2, subtract 40, multiply by 50, add the last two digits of your birth year, subtract 10500…. viola! Your birth date! Impress your friends! Now, use your birthday for some more math fun. My birthday is 11/01/1997 1) Create a quadratic function, f(x), in vertex form. The a should be between 4 and –4, the h will be your birth month, and the k will be your birth day. Write your equation below. 2) Using complete sentences, explain how to convert your birthday function into standard form 3) Graph your function. Include your graph below. 4) Using complete sentences, explain how to find the average rate of change for f(x) from x = 4 to x = 7. Need help desperately!

OpenStudy (anonymous):

Create a quadratic function, f(x), in vertex form. The a should be between 4 and –4, the h will be your birth month, and the k will be your birth day. Write your equation below. f(x)=3(x-11)2+1

OpenStudy (anonymous):

standard form is y(x) = a x^2 + b x + c where a, b, and c are numbers. the first step is "expand" (x-11)^2 = (x-11)(x-11) (you should have born in January.. you would have an easier problem)

OpenStudy (anonymous):

use FOIL (for example) to multiply it out. combine like terms. add in the +27 and you will have standard form

OpenStudy (anonymous):

The first step is to multiply (x-11)(x-11) the video will explain it faster (about 3 minutes) than we can do it by typing here...

OpenStudy (anonymous):

no one wants to read all these words did you get an answer yet?

OpenStudy (anonymous):

@satellite73 is it too much... :(

OpenStudy (anonymous):

too many words for me cut to the chase pick an \(a\), any \(a\) between \(-4\) and \(4\) make it easy

OpenStudy (anonymous):

I just wanted to make sure i explained it well to him. @satellite73

OpenStudy (anonymous):

i honestly do appreciate the help and you trying to explain it well for me to understand.. I just have little time and like i said I'm not too well with algebra..

OpenStudy (anonymous):

if any of you 2 can just quickly tell me how to do it? please..

OpenStudy (anonymous):

it would save both our times and get you your medals faster :-)

OpenStudy (anonymous):

@satellite73 doesn't need any more medals lol....

OpenStudy (anonymous):

@satellite73 mind helping? please